# Deep Learning Framework for History Matching CO2 Storage with 4D Seismic and Monitoring Well Data

### v1

@@ 6.2 History matching results
Figure 17: True model, three randomly sampled prior realizations, and three randomly sampled posterior realizations obtained using both monitoring well and interpreted seismic data. Geomodels shown in terms of $\log k_{x}$, with $k_{x}$ in md.
Although the posterior saturation fields shown in Figures 15 and 16 are at the coarse (interpreted seismic) scale, it is important to reiterate that our procedure provides high-resolution posterior geomodels. This is illustrated in Figure 17, where we display randomly sampled prior and posterior geomodel realizations ($\log k_{x}$ is the quantity displayed) along with the true model. The posterior models derive from history matching using both data types. It is evident that the three prior models show very different ranges for $\log k_{x}$, with notable scale variation between Prior 2 and Prior 3 (Figure 17(c) and (d)). The posterior realizations, by contrast, show similar ranges and variability in $\log k_{x}$, as well as general correspondence with the true model (Figure 17(a)). This is as expected since the posterior uncertainty in $\mu_{\log k}$ and $\sigma_{\log k}$ is relatively small.
@@ 7 Concluding remarks
In this work, we implemented a deep-learning-based framework for history matching of geological carbon storage operations when both 4D seismic and monitoring well data are available. Our emphasis was on early time predictions, which are particularly important because uncertainty is high and unexpected behavior must be quickly identified. Two different fit-for-purpose deep learning surrogate models were constructed – one for predicting interpreted (coarse-scale) 4D seismic saturation fields, and one for predicting monitoring well data. Both networks accept high-resolution geomodels as input. The network for interpreted 4D seismic involves a 3D U-Net architecture, while the network for monitoring well data applies a 1D U-Net. Both networks have multiple output channels, with each channel corresponding to a different time step. The two specialized networks are simpler to construct and less time consuming to train than a single overall network able to provide global high-fidelity predictions.
@@
The deep learning surrogate models were then applied for data assimilation using a hierarchical MCMC method. History matching results were constructed using only monitoring well data and using both monitoring well and interpreted 4D seismic data. The impact of 4D seismic, in terms of reducing posterior uncertainty in the metaparameters over that achieved using only monitoring well data, was clearly demonstrated. Predictions for CO2 plume shape and extent were also shown to be enhanced through inclusion of seismic data. A major advantage of our methodology is that the benefit of each data type can be quantified, and different strategies for combining highly-resolved local (monitoring well) data and coarse-resolution global (4D seismic) data can be quickly evaluated and thus optimized.

### v2

@@ 6.2 History matching results
Figure 20: True model, three randomly sampled prior realizations, and three randomly sampled posterior realizations obtained using monitoring well and interpreted saturation seismic data. Geomodels shown in terms of $\log k_{x}$, with $k_{x}$ in md.
Although the posterior saturation fields shown in Figures 18 and 19 are at the coarse (interpreted seismic) scale, it is important to reiterate that our procedure provides high-resolution posterior geomodels. This is illustrated in Figure 20, where we display randomly sampled prior and posterior geomodel realizations ($\log k_{x}$ is the quantity displayed) along with the true model. The posterior models derive from history matching using monitoring well and interpreted saturation seismic data. It is evident that the three prior models show very different ranges for $\log k_{x}$, with notable scale variation between Prior 2 and Prior 3 (Figure 20(c) and (d)). The posterior realizations, by contrast, show similar ranges and variability in $\log k_{x}$, as well as general correspondence with the true model (Figure 20(a)). This is as expected since the posterior uncertainty in $\mu_{\log k}$ and $\sigma_{\log k}$ is relatively small.
@@ 7 Concluding remarks
In this work, we implemented a deep-learning-based framework for history matching of geological carbon storage operations when both 4D seismic and monitoring well data are available. Our emphasis was on early time predictions, which are particularly important because uncertainty is high and unexpected behavior must be quickly identified, though our workflow is also applicable for longer time frames. Two different fit-for-purpose deep learning surrogate models were constructed – one for predicting interpreted (coarse-scale) 4D seismic saturation fields, and one for predicting monitoring well data. Both networks accept high-resolution geomodels as input. The network for interpreted 4D seismic involves a 3D U-Net architecture, while the network for monitoring well data applies a 1D U-Net. Both networks have multiple output channels, with each channel corresponding to a different time step. The two specialized networks are simpler to construct and less time consuming to train than a single overall network able to provide global high-fidelity predictions.
@@
The deep learning surrogate models were then applied for data assimilation using a hierarchical MCMC method. History matching results were constructed using only monitoring well data and using both monitoring well and interpreted 4D seismic data. The latter entailed either (binary) interpreted plume location data or interpreted saturation data. The impact of either type of 4D seismic data, in terms of reducing posterior uncertainty in the metaparameters over that achieved using only monitoring well data, was clearly demonstrated. Predictions for CO2 plume shape and extent were also shown to be enhanced through inclusion of seismic data. A major advantage of our methodology is that the benefit of each data type can be quantified, and different strategies for combining highly-resolved local (monitoring well) data and coarse-resolution global (4D seismic) data can be quickly evaluated and thus optimized.